OpenStudy (yacoub1993):

Find the slope of the tangent line to the graph of f at the given point. f(x) = x^2 + 5x at (4, 36)

OpenStudy (kc_kennylau):

That is the same as finding the derivative of the function at \(x=4\)

OpenStudy (yacoub1993):

how can i solve it

OpenStudy (kc_kennylau):

Firstly, can you find the derivative of \(f(x)\)? :)

OpenStudy (yacoub1993):

how?

OpenStudy (kc_kennylau):

Are you taught differentiation? :)

OpenStudy (yacoub1993):

no

OpenStudy (kc_kennylau):

@hartnn That's interesting

hartnn (hartnn):

you'll only need one formula of derivative here. \(\large \dfrac{d}{dx}x^n = n x^{n-1}\)

hartnn (hartnn):

so, can you find derivative of x^2 ? plug in n=2 in the formula

OpenStudy (yacoub1993):

@hartnn how can i find derivative of x^2 ?... Could you help me out trough

OpenStudy (kc_kennylau):

@hartnn Is there anyway to calculate tangent without calculus? :o

hartnn (hartnn):

i don't think so we can find the value of x for which the function is max or min, but slope of tangent.....we'll need derivative this question does belong to derivative topic right ? if not, then which topic ?limits?

OpenStudy (yacoub1993):

Yes the topic is derivative

hartnn (hartnn):

so you should have that formula which i gave you anyways, to find derivative of x^2 you'll just have to put n=2 in that formula what do u get ?

OpenStudy (yacoub1993):

and what does d stand for

OpenStudy (kc_kennylau):

\(\dfrac d{dx}\) is an expression meaning "differentiate"

hartnn (hartnn):

yes dx is the change in variable 'x' d f(x) is the change in function of x so, df(x)/dx is the change in function, with respect to change in x which we call as derivative of f(x)

OpenStudy (yacoub1993):

so what digit is d for... i am stuck

hartnn (hartnn):

\(\large \dfrac{d}{dx}x^n = n x^{n-1}\) put n=2 there what do u get ?

OpenStudy (yacoub1993):

\[\frac{ d }{ dx }4^{2}=2(4)^{2-1}\]

hartnn (hartnn):

why did u put x= 4 :O you'll only get d/dx (x^2) = 2x right ?

OpenStudy (yacoub1993):

yes

hartnn (hartnn):

using same formula can you tell me derivative of x ?

hartnn (hartnn):

hint : x means x^1, so put n=1 in that formula

OpenStudy (yacoub1993):

2x

OpenStudy (kc_kennylau):

try again :)

OpenStudy (yacoub1993):

x^2

hartnn (hartnn):

2x was the derivative of x^2 to get the derivative of 'x', plug in n=1 in that formula what u get if u put n=1 there?

OpenStudy (yacoub1993):

\[\frac{ d }{ dx }x ^{1} =1x ^{1-1}\]

hartnn (hartnn):

yes 1x^0 and whats x^0 =... ?

OpenStudy (yacoub1993):

0

hartnn (hartnn):

no...x^0 is just =1 so derivative of x just =1 got this?

OpenStudy (yacoub1993):

yup at last i got it

hartnn (hartnn):

good! so now try to find the derivative of x^2+5x ?

hartnn (hartnn):

Note : constants can be taken out of the derivative so d/dx (5x) = 5 d/dx (x)

OpenStudy (yacoub1993):

oka moment

OpenStudy (yacoub1993):

then

hartnn (hartnn):

could you find derivative of x^2+5x ? = derivative of x^2 + 5 derivative of x =. ...?

OpenStudy (yacoub1993):

0 right

hartnn (hartnn):

no how? we just found derivative of x^2 and derivative of x right ?

OpenStudy (yacoub1993):

sory derivative of x is 1

hartnn (hartnn):

yes, and derivative of x^2 ?

OpenStudy (yacoub1993):

its 1 again

hartnn (hartnn):

no... we found that derivative of x^2 =2x , right ? by putting n=2 in the formula

OpenStudy (yacoub1993):

ohh yes that right

hartnn (hartnn):

could you find derivative of x^2+5x ? = derivative of x^2 + 5 derivative of x =. ...?

OpenStudy (yacoub1993):

derivative of x^2+5x is 7x right

hartnn (hartnn):

why 7x ?

OpenStudy (yacoub1993):

2x + 5x = 7x

hartnn (hartnn):

2x is correct but not 5x 5 * derivative of x = 5*1 =5 got this?

OpenStudy (yacoub1993):

ohh

hartnn (hartnn):

because derivative of x is 1 !

OpenStudy (yacoub1993):

ohh ok

hartnn (hartnn):

so derivative of x^2+5x = 2x+5 right?

OpenStudy (yacoub1993):

yes

OpenStudy (yacoub1993):

so we go ahead and put the value of x and solve right

hartnn (hartnn):

so the slope of tangent at 4,36 is the derivative of graph at 4,36 so just put x=4 in your derivative yes, right!

OpenStudy (yacoub1993):

13 is the answer right

hartnn (hartnn):

absolutely :)

OpenStudy (yacoub1993):

woooo.. that was long steps for 1 question... U deserve more than medal @hartnn

OpenStudy (yacoub1993):

thnxs alot @hartnn

hartnn (hartnn):

don't worry for steps, once you have enough practice, you'll be able to solve derivatives questions mentally ;) you're welcome ^_^